# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ A.4 The proof of Lemma 1
Appendix B Index Set Equality
@@
If for all $\tau\in\mathcal{T}$, $c_{\tau}=0$ and $A$ is continuous defined on a cone with $Im(A)=\mathbb{R}^{+}$ such that $A(0)=0$, $\nabla A(0)=0$ and $\lim_{\|\mu\|\to\infty}\frac{A(\mu)}{\|\mu\|}=+\infty$ we have $\mathcal{F}_{\mathcal{T}}=\mathcal{G}_{\mathcal{T}}$.
@@
Consider $\tau$ in $\mathcal{G}_{\mathcal{T}}$. For all $\tau^{\prime}\neq\tau$, the inverse image of the open set $(0,+\infty)$ by linear functions $g_{\tau}-g_{\tau^{\prime}}$ is a non-empty intersection of half-spaces. Moreover, there exists $(\lambda_{0},\mu_{0})\in Im(A)\times\mathcal{D}_{A}$ such that for all $\tau^{\prime}\neq\tau$ in $\mathcal{T}$ we have: $g_{\tau}(\lambda_{0},\mu_{0})>g_{\tau^{\prime}}(\lambda_{0},\mu_{0})$. Using condition $c_{\tau}=c_{\tau^{\prime}}=0$ and $\alpha>0$ we get by linearity
@@
It remains to prove that the proposed assumptions over $A$ are sufficient to find such an alpha proving the set equality. We consider the same half-line $\{\alpha(\lambda_{0},\mu_{0}),\alpha>0\}$. By linear approximation in 00, we have $A(\epsilon)=A(0)+\langle\epsilon,\nabla A(0)\rangle+o(\|\epsilon\|)=o(\|\epsilon\|)$ and locally, we get $A(\alpha\mu_{0})=o(\|\alpha\mu_{0}\|)<\alpha\lambda_{0}$ for small $\alpha$. Notice that as $Im(A)=\mathbb{R}^{+}$ and the cone is open, we have $\lambda_{0}>0$ and can choose $\|\mu_{0}\|>0$. The half-line is therefore above the manifold ("$(\alpha\lambda_{0},\alpha\mu_{0})>(A(\alpha\mu_{0}),\alpha\mu_{0})$").
@@
Appendix C Stirling numbers of the first kind in terms of the harmonic series
To obtain numerical values of $U^{p}_{n}$ and $V_{n}^{p}$ of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\dots,n-1\}}$ for $1\leq p\leq 5$ we need to know the expressions of $\begin{bmatrix}n\\ m\\ \end{bmatrix}$ for $m=0,\dots,6$. As shown in [58], the general formula for the Stirling numbers of the first kind in terms of the harmonic series is
@@ Appendix C Stirling numbers of the first kind in terms of the harmonic series
where $\sigma_{k+1}=\sum_{i=1}^{n-1}\frac{1}{i^{k+1}}$ is the partial sums of the harmonic series $\left\{\frac{1}{i^{k+1}}\right\}_{i=1,2,\dots}$ and $(1-m)_{k}$ are the Pochhammer symbols. Applying the formula (14) for $m=0,\dots,6$, we get
@@
Appendix D Detection of multiple changes in the parameter of a distribution in the exponential family
We define the set $\mathcal{M}_{n}^{K}$ of all possible segmentations in $K$ changepoints of time series $\{y_{t}\}_{t=1,\dots,n}$ as:
$\mathcal{M}_{n}^{K}=\{\tau=(\tau_{0},\tau_{1},\dots,\tau_{K},\tau_{K+1})\in\mathbb{N}^{K+2}|0=\tau_{0}<\tau_{1}<\dots<\tau_{K}<\tau_{K+1}=n\}.$
For any segmentation $\tau$ in $M_{n}^{K}$ we define its size as $|\tau|=K$. We denote $\mathcal{M}_{n}^{\infty}$ as the set of all possible segmentations of $\{y_{t}\}_{t=1,\dots,n}$:
$\mathcal{M}_{n}^{\infty}=\bigcup_{K}M_{n}^{K}\,,$ (15)
and take the convention that $M_{n}^{\infty-1}=M_{n}^{\infty}\,$.
From a statistical perspective given a subset of all segmentations $\mathcal{M}_{n}^{K}$ for $K$ in ${}\mathbb{N}\cup\{\infty\}$ it makes sense to recover the segmentation maximising a penalised likelihood of our (transformed as in Section 2.2) data $\{x_{t}\}_{t=1,\dots,n}$:
$C_{n}^{K}=\max_{\tau\in\mathcal{M}_{n}^{K}}\left\{\sum_{k=1}^{|\tau|+1}\max_{\eta_{k}}\left[\sum_{t=\tau_{k-1}+1}^{\tau_{k}}2\log(f_{X}(x_{t}|\eta_{k}))-\beta\right]\right\}\,,$ (16)
where $\beta$ is a penalty value in $\mathbb{R}^{+}$.
The previous equation encompasses two common multiple changepoints strategies. In the first, which we call Segment Neighbourhood (SN) [22] the user is given a maximum number of changepoints $K_{max}$ and solve Equation (16) for all $K\leq K_{max}$ (setting $\beta$ to 00). In the second step all these solutions are compared using an appropriate model selection criteria such as [28, 73] to select the number of changepoints. In the second strategy, which we call Optimal Partitioning (OP) [25] the position and number of changepoints are estimated in one go solving (16) for $K=\infty$ and picking an appropriate value for $\beta$ such as $2\hat{\sigma}^{2}\log(n)$ in a Gaussian setting with $p=1$ [74] where $\hat{\sigma}^{2}$ is an estimate of the Gaussian variance.
Finding the maximum of (16) in either case can be done using dynamic programming. For any segmentation $\tau$ with $|\tau|>1$, observe that the first $|\tau|$ terms in the previous sum are by definition smaller than $C_{\tau_{|\tau|}}^{K-1}$. Hence, we only need to consider the position of the last changepoint and we recover the update rule of the SN [22] and OP [25] algorithms:
$C_{t}^{K}=\max_{i<t}\left\{C_{i}^{K-1}+\max_{\eta}\left[\sum_{s=i+1}^{t}2\log(f_{X}(x_{s}|\eta))-\beta\right]\right\}\ .$ (17)
With this recursion solved for increasing $t$, and, in the case of SN, increasing $K$.
Directly solving this recursion has a computational cost that is quadratic in $n$. This has led to pruning ideas that speed up the computation whilst still solving the maximisation problem exactly. The idea of these pruning rules is that at any time $t$ we need only consider a subset of possible changepoint locations. The most efficient pruning rules are based on so-called functional pruning (as in the pDPA or FPOP algorithms [32, 33]). The idea is to consider the best segmentation of the data to time $t$ as a function of the parameter of the final segment. For each value of this parameter there will be an optimal position of the most recent changepoint prior to $t$, and at the next iteration, we need only consider these positions and $t$ for the possible position of the most recent changepoint prior to $t+1$. Implementing this algorithm requires calculating regions of the parameter space that are optimal for each value of the most recent changepoint. When $p=1$ this is straightforward as we are manipulating intervals, but for $p>1$ this becomes challenging, and often computationally prohibitive, as it involves computing the intersection and subtraction of many convex sets in $\mathbb{R}^{p}$ [35, 36].
D.1 Connection between the Detection of Multiple changepoints and the half-space intersection problem
In Section 4, we established that the maximum likelihood estimator of a single changepoint problem corresponds to a vertex on the hull of a particular set of points. This result is primarily based on Theorem 4, which follows from the Half-space Intersection Theorem 3.
In the following theorem, we establish that the last changepoint of the maximum likelihood segmentation (see Equation (16)) of a multiple changepoint problem corresponds to a supporting hyperplane of a particular half-space intersection problem defined as in Theorem 3.
Theorem 8.
The last changepoint $\hat{\imath}$ of the segmentation $\hat{\tau}$ maximising the likelihood (16) is such that $H_{\hat{\imath}}$ is a supporting hyperplane of $\cap_{i}H_{i}$ where $i$ is in $\{1,\ldots,n-1\}$ and $H_{i}$ is defined as the hyperplane
$H_{i}=\{\kappa,\lambda,\mu\ |\ h_{i}(\kappa,\lambda,\mu)=a_{i}\lambda-2\langle b_{i},\mu\rangle+c_{i}-\kappa\leq 0\},$
with $a_{i}=i$, $b_{i}=\sum_{t=1}^{i}x_{t}$ and $c_{i}=C_{i}^{K-1}-\sum_{t=1}^{i}B^{\prime}(x_{t})$.
Proof.
We start from Equation (16). There is a value $\hat{\eta}$ for which $\hat{\tau}$ is optimal. Now if we subtract $\sum_{t=1}^{n}\Big{\{}2\log(f_{X}(x_{t}|\eta))-\beta\Big{\}}$, $\hat{\eta}$ is still optimal for $\eta=\hat{\eta}$. Now applying Theorem 3 and restricting our attention to the last changepoint as explained for Equation (9) we get the desired result. ∎
Contrary to the single changepoint case, we do not propose an algorithm based on this connection. Nonetheless, in a manner similar to the MdFOCuS algorithm, we think one could restrict the set of last changepoints to consider in the standard SN or OP algorithm using the output of a half-space intersection algorithm.
Appendix E Additional Empirical evaluations
E.1 Run time as a function of a slope parameter
Now we consider the empirical run times of the MdFOCuS algorithm with a known pre-change parameter $\eta_{1}$ as a function of a slope parameter $\alpha$. Using the implementation of Algorithm 1 for a change in the mean of multi-dimensional Gaussian signal (see Table 1) in R/C++ using the qhull library (see link in the Supplementary Material) we studied how the $\alpha$ parameter affects its running time. To do this, we generate $100$ time series with $n=10^{5}$ data points (without changepoints, and $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d.) in dimension $p=1,2$ and $3$. In Figure 11 we report the average run time as a function of $\alpha$ for MdFOCuS0 (that is for a known value of $\eta_{1}$) and values of $\alpha$ in $\left\{1+\{2^{i}\}_{i\in\{-4,\dots,0\}},3,\{2^{j}\}_{j\in\{2,\dots,7\}}\right\}$. The value of $maxSize$ was always initialised to $p+2$ and $\beta$ was always set to $1$. We observe a minimum run time for $\alpha$ in $[2,4]$. As a consequence in the rest of the simulations we always set $\alpha$ to $2$.
Figure 11: Run time of MdFOCuS0 as a function of $\alpha$ for $p$-variate time series with $n=10^{5}$ data points in dimension $p=1,2$ and $3$. We simulated $100$ i.i.d. Gaussian data $\mathcal{N}_{p}(0,I_{p})$ and report the average run time.
E.2 Multidimensional Poisson Model
We simulated time series with a Poisson distribution : $x_{1:n}$ ($x_{t}\sim\mathcal{P}(\mathbf{1}_{p})$, $t=1,\dots,n$). We varied $p$ and $n$ as in Section 4.2.
E.2.1 Empirical validation of the bound
@@
Using the same procedure as in Section 4.2, we evaluate the number of faces and vertices of the convex hull $\{P(\tau)\}_{\tau\in\{1\dots,n-1\}}$. The results are presented in Figure 13. We note that the results are similar to the Gaussian case (see Figure 4).
E.2.2 Empirical run times of the algorithm as a function of dimension
@@
Using the same procedure as in Section 4.3 we evaluate the run times of MdFOCuS0 and MdFOCuS for Poisson distribution. The results are presented in Figure 13. We note that the results are similar to the Gaussian case (see Figure 4).
@@
Figure 12: The number of faces (left) and vertices (right) of the convex hull built from the set of points $\{P(i)\}_{i\in\{1\dots,n-1\}}$ for dimension $1\leq p\leq 5$. We simulated $100$ i.i.d. $p$-variate Poisson data $\mathcal{P}(\textbf{1}_{p})$ for each $n$ from $(2^{11}+1)$ to $(2^{23}+1)$. Dashed lines correspond to the expected number of faces (left) and vertices (right) presented in Theorem 6. We consider a maximum running time of 40 minutes for the Quickhull algorithm. That is why we don’t have data points for $p=4$ and $5$ for large number of observations.
@@
Figure 13: Run times in seconds of MdFOCuS with known (dashed line) and unknown(full line) pre-change parameter $\eta_{1}$, both with $\alpha=2$, $\beta=1$ and $maxSize=14$, in dimension $p=1,\dots,6$ using time series $x_{1:n}$ (without change, $x_{t}\sim\mathcal{P}(\textbf{1}_{p})$ i.i.d.). Run times are averaged over $100$ data sets. We considered a maximum running time of $20$ minutes (horizontal dotted black line) for the algorithms. That is why we do not report run times for $p=4,5$ and $6$ for large number of observations. We report here the slope of the curve obtained using a simple linear regression: for known $\eta_{1}$: $\approx 1.013$ ($p=1$), $\approx 1.126$ ($p=2$), $\approx 1.192$ ($p=3$), $\approx 1.210$ ($p=4$), $\approx 1.172$ ($p=5$) and $\approx 1.250$ ($p=6$); for unknown $\eta_{1}$: $\approx 1.018$ ($p=1$), $\approx 1.138$ ($p=2$), $\approx 1.213$ ($p=3$), $\approx 1.240$ ($p=4$), $\approx 1.190$ ($p=5$) and $\approx 1.254$ ($p=6$).
E.3 50-dimensional time series, controlled by false positive rate
@@
We present an additional simulation study to compare the 1d approximation proposed by [34] (here called FOCuS), with the high-dimensional approximation for MdFOCuS (in Table 5). We compare MdFOCuS 2d Sp, which utilizes a two-dimensional approximation of the convex hull alongside the partial likelihood ratio statistic (see subsection G). To facilitate a fair comparison, we tune the thresholds of all methods to achieve the same pre-change false positive rate of 0.01. We applied a Bonferroni correction to cope with multiple thresholds. The simulation scenarios mirror those of the main study but with the changepoint occurring at time 10000. This is to emphasize the impact of the approximations. We expect the two-dimensional approximation to outperform the one-dimensional approximation proposed by [34], particularly for small-magnitude changes. However, as observed in the main study, the original FOCuS method is still superior when observing a change in one dimension, as, despite in that case the statistics being equal, its threshold is slightly lower (because it considers less statistics in total).
@@
Table 5: Average detection delay with 50-dimensional time series for a change at time $t=10000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold. Results are by controlling the false positive rate.
@@
magnitude density FOCuS0 MdFOCuS0 2d Sp FOCuS0 (est 500) FOCuS MdFOCuS 2d Sp
@@
E.4 Additional Plug-in estimated methods

### v3

@@ B.4 Proof of Theorem 6
Based on Lemma 2 properties $(Ex)$ and $(GP)$ hold and we can apply formula (2) from Kabluchko et al. [2017] on the expected values of $U_{n}^{p}$ and $V_{n}^{p}$. This completes the proof of Theorem 6.
Remark 2.
Theorem 6 gives the expectation for i.i.d. observations. One might wonder what happens if there is a changepoint in the distribution at $i<n$. Note that the points on the hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ are necessarily on the hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,i\}}$ or on the hull of $\{P(\tau)\}_{\tau\in\{i+1,\ldots,n-1\}}$. Thus applying the theorem on all points from $1$ to $i$ and from $i+1$ to $n-1$ we expect at most $\mathbb{E}(V_{i+1}^{p})+\mathbb{E}(V_{n-i}^{p})\leq 2\mathbb{E}(V_{n}^{p})$ points on the hull (as the expectation of $V_{i}^{p}$ is increasing with $i$). Thus if there is a changepoint (and assuming we have not stopped because the GLR statistics is lower than the specified threshold) the expected number of candidate changepoints is still in $\mathcal{O}(\log(n)^{p})$.
B.5 Empirical validation of the bound of Theorem 6
We empirically tested the validity of Theorem 6 for $p\leq 5$. We use the Gaussian distribution as an example of a distribution with continuous density and independent dimensions. To be specific, for $p$ from $1$ to $5$, we simulated time series $x_{1:n}$, with $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d. We varied the time series length from $2^{10}+1(\approx 10^{3})$ to $2^{23}+1(\approx 8\times 10^{6})$ and simulated $100$ data sets for each length. We used the QuickHull algorithm Barber et al. [1996] implemented in the R package geometry Habel et al. [2023] to evaluate the number of faces and vertices of the convex hull of $\{P((\tau)\}_{\tau\in\{1,\ldots,n-1\}}$. It can be seen in Figure 10 that the observed number of faces and vertices are close to their theoretical expectations (from Theorem 6). Additional details on the calculations of the Stirling numbers are provided in Supplementary Material E.
The proof of Theorem 6 depends on the fact that the distribution is continuous. We thus empirically tested the robustness of the expectations to this assumption by considering time series drawn with a discrete i.i.d. Poisson distribution. The results closely resemble the Gaussian case and details can be found in Section F.1.
Appendix C Exponential Family Models
Distribution $\eta:=r(\theta)$ $x:=s(y)$ $A^{\prime}(\eta)$ $B^{\prime}(x)$
Gaussian (changes in mean) $\theta$ $y$ $\|\eta\|^{2}$ $\|x\|^{2}+\log(2\pi)^{p}$
Binomial $\log\frac{\theta}{1-\theta}$ $y$ $2m\langle\log(1+e^{\eta}),\mathbf{1}_{p}\rangle$ $-2\left\langle\log\begin{pmatrix}m\\ x\end{pmatrix},\mathbf{1}_{p}\right\rangle$
Exponential $-\theta$ $y$ $-2\langle\log(-\eta),\mathbf{1}_{p}\rangle$ 0
Examples of models and change-types covered by our setting are given in Table 4. Our algorithms can also deal with data where different components of a multivariate stream are from different models.
Our setting also covers the change in mean ($\theta_{1}$) and variance ($\theta_{2}$) (so $p^{\prime\prime}=2$) of a Gaussian signal in $\mathbb{R}$ (so $p^{\prime}=1$). In this slightly more complex case we have $p^{\prime}\neq p=2$ and we can define $s$, $r$, $A^{\prime}$ and $B^{\prime}$ as follows
$\theta=(\theta_{1},\theta_{2}),~{}\eta=\left(\frac{\theta_{1}}{\theta_{2}},-\frac{1}{2\theta_{2}}\right),~{}s(y)=(y,y^{2}),~{}A^{\prime}(\eta)=\frac{\theta_{1}^{2}}{\theta_{2}}-\log(\theta_{2}),~{}B^{\prime}(x)=\log(2\pi).$
Appendix D Index Set Equality
@@
If $A$ is continuous and defined on a cone with $Im(A)=\mathbb{R}^{+}$ such that $A(0)=0$, $\nabla A(0)=0$ and $\lim_{\|\mu\|\to\infty}\frac{A(\mu)}{\|\mu\|}=+\infty$ we have $\mathcal{F}_{\mathcal{T}}=\mathcal{G}_{\mathcal{T}}$.
@@
Consider $\tau$ in $\mathcal{G}_{\mathcal{T}}$. For all $\tau^{\prime}\neq\tau$, the inverse image of the open set $(0,+\infty)$ by linear functions $g_{\tau}-g_{\tau^{\prime}}$ is a non-empty intersection of half-spaces. Moreover, there exists $(\lambda_{0},\mu_{0})\in Im(A)\times\mathcal{D}_{A}$ such that for all $\tau^{\prime}\neq\tau$ in $\mathcal{T}$ we have: $g_{\tau}(\lambda_{0},\mu_{0})>g_{\tau^{\prime}}(\lambda_{0},\mu_{0})$. For $\alpha>0$ we get by linearity
@@
It remains to prove that the proposed assumptions over $A$ are sufficient to find such an alpha proving the set equality. We consider the same half-line $\{\alpha(\lambda_{0},\mu_{0}),\alpha>0\}$. By a linear approximation about 00, we have $A(\epsilon)=A(0)+\langle\epsilon,\nabla A(0)\rangle+o(\|\epsilon\|)=o(\|\epsilon\|)$ and locally, we get $A(\alpha\mu_{0})=o(\|\alpha\mu_{0}\|)<\alpha\lambda_{0}$ for small $\alpha$. Notice that as $Im(A)=\mathbb{R}^{+}$ and the cone is open, we have $\lambda_{0}>0$ and can choose $\|\mu_{0}\|>0$. The half-line is therefore above the manifold ("$(\alpha\lambda_{0},\alpha\mu_{0})>(A(\alpha\mu_{0}),\alpha\mu_{0})$").
@@
Appendix E Stirling numbers of the first kind in terms of the harmonic series
To obtain numerical values of $U^{p}_{n}$ and $V_{n}^{p}$ of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\dots,n-1\}}$ for $1\leq p\leq 5$ we need to know the expressions of $\begin{bmatrix}n\\ m\\ \end{bmatrix}$ for $m=0,\dots,6$. As shown in Adamchik [1997], the general formula for the Stirling numbers of the first kind in terms of the harmonic series is
@@ Appendix E Stirling numbers of the first kind in terms of the harmonic series
where $\sigma_{k+1}=\sum_{i=1}^{n-1}\frac{1}{i^{k+1}}$ is the partial sums of the harmonic series $\left\{\frac{1}{i^{k+1}}\right\}_{i=1,2,\dots}$ and $(1-m)_{k}$ are the Pochhammer symbols. Applying (14) for $m=0,\dots,6$, we get
@@
Appendix F Additional Empirical evaluations
F.1 Multidimensional Poisson Model
@@
We simulated time series with each data-point simulated a $p$ independent Poisson random variables with mean 1. We varied $p$ and $n$ as in Section B.5. Using the same procedure as in Section B.5, we evaluate the number of faces and vertices of the convex hull $\{P(\tau)\}_{\tau\in\{1\dots,n-1\}}$. The results are presented in Figure 11. We note that the results are similar to the Gaussian case (see Figure 10).
@@
Next, using the same procedure as in Section 1, we evaluated the run times of MdFOCuS for a known or unknown pre-change parameter for Poisson distribution. The results are presented in Figure 12. We note that the results are similar to the Gaussian case (see Figure 3).
@@
Figure 11: The number of faces (left) and vertices (right) of the convex hull built from the set of points $\{P(i)\}_{i\in\{1\dots,n-1\}}$ for dimension $1\leq p\leq 5$. Results averaged over 100 realisations of independent $p$-dimensional Poisson random variables with mean 1 data. Dashed lines correspond to the expected number of faces (left) and vertices (right) presented in Theorem 6. We consider a maximum running time of 40 minutes for the Quickhull algorithm, hence no data points for $p=4$ and $5$ for large numbers of observations.
@@
Figure 12: Run times in seconds of MdFOCuS with known (dashed line) and unknown (full line) pre-change parameter $\eta_{1}$, both with $\alpha=2$, $\beta=1$ and $maxSize=14$, in dimension $p=1,\dots,6$ using time series $x_{1:n}$ simulated from the independent Poisson model. Run times are averaged over $100$ data sets. We considered a maximum running time of $20$ minutes (horizontal dotted black line) for the algorithms, hence no run times for $p=4,5$ and $6$ for large numbers of observations. The slope of the curves was obtained using a simple linear regression: for known $\eta_{1}$: $\approx 1.013$ ($p=1$), $\approx 1.126$ ($p=2$), $\approx 1.192$ ($p=3$), $\approx 1.210$ ($p=4$), $\approx 1.172$ ($p=5$) and $\approx 1.250$ ($p=6$); for unknown $\eta_{1}$: $\approx 1.018$ ($p=1$), $\approx 1.138$ ($p=2$), $\approx 1.213$ ($p=3$), $\approx 1.240$ ($p=4$), $\approx 1.190$ ($p=5$) and $\approx 1.254$ ($p=6$).
F.2 Run time as a function of a slope parameter
We consider the empirical run times of the MdFOCuS algorithm with a known pre-change parameter $\eta_{1}$ as a function of parameter $\alpha$. Using the implementation of Algorithm 1 for a change in the mean of multi-dimensional Gaussian signal (see Table 4) in R/C++ using the qhull library we studied how the $\alpha$ parameter affects its run time. To do this, we generate $100$ time series with $n=10^{5}$ data points, with each data point an independent realisation of a standard $p$-dimensional Gaussian random variable, in dimension $p=1,2$ and $3$. In Figure 13 we report the average run time as a function of $\alpha$ for MdFOCuS for the pre-change parameter known case. The value of maxSize was always initialised to $p+2$ and $\beta$ was always set to $1$. We observe a minimum run time for $\alpha$ in $[2,4]$. As a consequence, in the rest of the simulations we always set $\alpha$ to $2$.
Figure 13: Run time of MdFOCuS as a function of $\alpha$ for $p$-variate time series with $n=10^{5}$ data points in dimension $p=1,2$ and $3$. Results averaged over 100 simulations.
F.3 50-dimensional time series, controlled by false positive rate
@@ F.2 Run time as a function of a slope parameter
We present, in Table 5, an additional simulation study to compare the simple one-dimentional multivariate implementation proposed by Romano et al. [2023] (here called FOCuS), with the high-dimensional approximation for MdFOCuS using the two-dimensional approximation of the convex hull alongside the partial likelihood ratio statistic (see Section 4.5 and subsection H). To facilitate a fair comparison, we tune the thresholds of all methods to achieve the same pre-change false positive rate of 0.01. We applied a Bonferroni correction to cope with multiple thresholds. The simulation scenarios mirror those of the main study but with the changepoint occurring at time 10000. This is to emphasize the impact of the approximations. We expect the two-dimensional approximation to outperform the one-dimensional approximation proposed by Romano et al. [2023], particularly for small-magnitude changes. However, as observed in the main study, the original FOCuS method is still superior when observing a change in one dimension, as, despite in that case the statistics being equal, its threshold is slightly lower (because it considers fewer statistics in total).
@@
Table 5: Average detection delay with 50-dimensional time series for a change at time $t=10000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold. Results are by controlling the false positive rate. FOCuS (est 500) estimates the pre-change parameter from training data of length 500.
@@
magnitude sparsity FOCuS MdFOCuS FOCuS (est 500) FOCuS MdFOCuS
@@
F.4 Additional Plug-in estimated methods